OpenStudy (anonymous):

if g(x)=3x^2-x+17 then g(2) 21 24 27 29 31

OpenStudy (anonymous):

27

OpenStudy (hoblos):

put 2 in the place of x... g(2) = 3(2)^2 -2 + 17 = 3(4) +15 = 12+15 =27